Assume that 0.401 g of diborane is combusted in a calorimeter whose heat capacity (Ccalorimeter) is 7.854 kJ/°C at 25.59°C. What is the final temperature of the calorimeter?
1
Expert's answer
2020-11-04T14:24:37-0500

=> Heat of combustion of diborane = -1941KJ/mol.

=> for 0.401 g Diborane , mole = 0.0154

So heat of combustion = 0.0154"\times (-1941) = -29.89 KJ" .

Now if final temperature is T, then from equation "-29.89KJ = C(T-25.59)"

"=> T-25.59 = -3.805"

=> Final temperature = "T= 21.78^0 C"